OpenStudy (highschoolmom2010):

Use what you know about trigonometric ratios to show that this equation is an identity.

OpenStudy (highschoolmom2010):

Sin (x)=cos (x) * tan (x)

OpenStudy (highschoolmom2010):

doesnt make sense to me

OpenStudy (anonymous):

you know that tanx = sinx/ cosx/

ganeshie8 (ganeshie8):

substitute the sin cos tan wid their ratios (opp, adj, hyp)

OpenStudy (anonymous):

so sub tanx= sinx/ cosx. this yields left side. left side= right side. QED

OpenStudy (anonymous):

tanx= sinx/ cos x because tanx by definition is opp/ adj. sinx= opp/ hyp cos= adj/hyp.

OpenStudy (highschoolmom2010):

that confused me waay more ^^

OpenStudy (anonymous):

just a note you can use frac{a}{b} in equations to make fractions ex. \[\frac{a}{b}\]

ganeshie8 (ganeshie8):

work left side separately work right side separately show both are equal

OpenStudy (anonymous):

\[\tan x = \frac{\sin x}{\cos x}\]

ganeshie8 (ganeshie8):

left side :- sin x = opp/hyp ---------------(1)

OpenStudy (anonymous):

Okay, i will take it really slowly.

OpenStudy (anonymous):

sinx= opp/ hyp= left side.

ganeshie8 (ganeshie8):

take the right side, substitute opp/adj/hyp, simplify and see if u can get the same as left side

OpenStudy (anonymous):

cosx= adj/ hyp; tan x= opp/adj . cosx* tanx= (adj/hyp) (opp/adj); notice adj can be cancelled out--> opp/ hyp= sinx. Notice it is same as sinx= left side.

OpenStudy (highschoolmom2010):

\[\frac{ adj }{ hyp }*\frac{ opp }{ adj}\] im really confused :/

ganeshie8 (ganeshie8):

keep going, one more step. confusion wil go away

OpenStudy (phi):

adj/adj = 1

OpenStudy (anonymous):

top adj can be cancelled out with bottom adj, and you are left with opp/hyp= sinx

OpenStudy (highschoolmom2010):

ok so i was almost there \[\sin (x)=\frac{ opp }{ hyp }\]

OpenStudy (highschoolmom2010):

like that

OpenStudy (highschoolmom2010):

brb yall meh baby is hungry

OpenStudy (phi):

the other way is to say \[ \tan(x) = \frac{\sin(x)}{\cos(x)}\] replace tan(x) in \[ \sin(x)= \cos(x) \tan(x) \\ \sin(x)= \cos(x)\frac{\sin(x)}{\cos(x)}\] then notice that cos(x)/cos(x) is 1 you get \[ \sin(x)=\sin(x) \]

ganeshie8 (ganeshie8):

Right hand side :- \(\large \frac{ adj }{ hyp }*\frac{ opp }{ adj} \) \(\large \frac{ \cancel{adj} }{ hyp }*\frac{ opp }{ \cancel{adj}} \) \(\large \frac{ opp }{ hyp } \) ------------(2)

ganeshie8 (ganeshie8):

from (1) and (2), left hand side = right hand side. so its an identity.

OpenStudy (highschoolmom2010):

oh ok so how do i write it into words

ganeshie8 (ganeshie8):

you wanto write math into words ? hmm

ganeshie8 (ganeshie8):

just showing the proof wont do ha ?

OpenStudy (highschoolmom2010):

no i have to write it into words :(

ganeshie8 (ganeshie8):

okay then describe wat we just did, you may start like below :- since left hand side is sinx, and we knw that sinx = opp/hyp, proving right hand side simplifies to opp/hyp is sufficient to prove that given equation is an identity. right hand side we have cosx * tanx. substituting their respective trig ratios, we get adj/hyp * opp/adj....

OpenStudy (highschoolmom2010):

ty

ganeshie8 (ganeshie8):

np, im sure you can conclude :)

OpenStudy (highschoolmom2010):

yea i think i can :D

OpenStudy (highschoolmom2010):

@phi @katherine.ok @ganeshie8 thank you all for for help